You can park in a a lot for 15$ an hour for at most 4 hours. A partial hour rounds to the nearest whole hour. find the function, domain and range.
I thought the function was y=15x and the domain was all real numbers.

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

your equation is y = 15 * x for x <= 4.
you domain is 0 <= x <= 4
your range is 0 <= 15*4 = 60

the function is y = 15, but the domain (the possible values of x, is from 0 to 4.
the range is the possible values of y for all possible values of x, which is from 0 to 60.

here's a graph.



y can't be less than 0 because x can't be less than 0.

y can't be greater than 60 because x can't be greater than 4.


Answer by ikleyn(52795)

.
You can park in a a lot for 15$ an hour for at most 4 hours. A partial hour rounds to the nearest whole hour.
find the function, domain and range.
I thought the function was y=15x and the domain was all real numbers.
~~~~~~~~~~~~~~


This post has a problem.

From the formal point of view, the cost function is not defined completely/properly in the post.

Namely,  for  EACH  of the values of  t= 0.5  hours,  1.5  hours,  2.5  hours and  3.5  hours,  there are
TWO  closest integers,  and the post does not define properly/accurately,  which one of the two closest integer values to select.

So,  there is a  HUGE  HOLE  in the problem formulation, which remove the posted text from the class of correctly posed  Math problems.


If you will edit and re-post your problem,  I will show you the correct solution.

Notice that the answer by tutor  @Theo is not correct,  unfortunately.

The cost function in this problem is  DIFFERENT  from that shown in his post.